OpenStudy (anonymous):

Can anyone help me with this problem please? 3x-2+X-4+3-2=0 I need a walk through thanks!

Vocaloid (vocaloid):

hint: what is 3x + x?

Vocaloid (vocaloid):

that's only the first step... now, what is -2-4+3-2?

Vocaloid (vocaloid):

still there?

Vocaloid (vocaloid):

1. start by combining all the x's 2. add all the other numbers together

Vocaloid (vocaloid):

-2-4+3-2 = ? use a calculator if you are unsure

Vocaloid (vocaloid):

that's what i'm trying to do

Vocaloid (vocaloid):

you did the first step very well, 3x + x = 4x

Vocaloid (vocaloid):

now the second step is to add everything else -2-4+3-2 = ?? you can use a calculator if you want

Vocaloid (vocaloid):

you are allowed to say "I'm not sure"

Vocaloid (vocaloid):

so.. did you solve it or do you still need help?